OpenStudy (arbyj):

Someone me help solve this; 4a+3=60a-45 I have the answer but I want to make sure its correct (Please show work)

OpenStudy (anonymous):

so first, get all of the same term on one side: 4a + 3 = 60a - 45 48 = 56a <<< then just divide by 56 on both sides a = 48/56 <<<then just simplify

OpenStudy (arbyj):

the answer would be .86, right? @uakid94

OpenStudy (anonymous):

Sorry for the late response but yes... you are correct. Well done. :)